Lucifer February Com Night 4 February 9-10 2009

Observers: N.Ageorges, W.Seifert, V.Knierem (LUCIFER), DThompson
Telescope Operator: W.Wack
SW Support: PGrenz on call because of bad weather
Telescope Support:


Never opened. Snow after sunset.

Mount oscillation problems, see IT#2002. Traced to HBS oil that was too cold.

-- DavidThompson - 11 Feb 2009
Topic revision: r1 - 11 Feb 2009, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback